OpenStudy (anonymous):

2r^2-πr^2 factored form

OpenStudy (zzr0ck3r):

\(2r^2-\pi r^2=r^2(2-\pi)\)

OpenStudy (zzr0ck3r):

all you can do is factor out the \(r^2\)

OpenStudy (anonymous):

thank you